The ratio of grey ties to black ties on display is 5 to 2. Use the ratio table to determine how many black ties will be on display if there are 35 gray ties


1
Expert's answer
2021-12-14T19:47:00-0500

Let "x" black ties will be on display if there are 35 gray ties. Then


"\dfrac{5}{2}=\dfrac{35}{x}"

"x=\dfrac{2(35)}{5}"

"x=14"

"14" black ties will be on display if there are 35 gray ties.